# A delayed interior area-to-height estimate for the Curve Shortening Flow

### v1

@@ 5.1 Constructing the flow
We start with a sequence $(u^{n}_{0})_{n\in\mathbb{N}}\subset\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R})$ of smooth and compactly supported functions for which $u^{n}_{0}\mathscr{L}^{1}\rightharpoonup\nu\in\mathcal{M}(\mathbb{R})$ and $u^{n}_{0}\to u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\Omega)$, and with another sequence $(U^{n}_{0})_{n\in\mathbb{N}}\subset\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R})$ which converges in the appropriate senses to $|\nu|\in\mathcal{M}_{*}^{+}(\mathbb{R})$ and satisfies $|u^{n}_{0}|\leq U^{n}_{0}$ on $\mathbb{R}$ for each $n\in\mathbb{N}$. We can construct such sequences by cutting-off and mollifying the positive and negative parts of $\nu\in\mathcal{M}_{*}(\mathbb{R})$ simultaneously, and then recombining the mollifications (see e.g. [Rud91, Ch. 6]).
Let $u^{n}\in\mathrm{C}^{\infty}(\mathbb{R}\times[0,\infty)\!)$ be the Ecker–Huisken flow of $u^{n}_{0}$, and let $U^{n}\in\mathrm{C}^{\infty}(\mathbb{R}\times[0,\infty);\mathbb{R}_{\geq 0})$ be the Ecker–Huisken flow of $U^{n}_{0}$. The Avoidance Principle (Theorem A.7) applied to ±$\pm u^{n}$ and $U^{n}+\eta$ for $\eta\searrow 0$ shows that $|u^{n}(t)|\leq U^{n}(t)$ on $\mathbb{R}$ for all $t\in[0,\infty)$.
We wish to take a locally smooth limit of the flows $u^{n}$ (and, for technical reasons in §5.2, the flows $U^{n}$) on ×$\mathbb{R}\times(0,\infty)$. By the Arzelà–Ascoli Theorem, it suffices to prove $\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$–bounds on the sequences of GCSFs which are independent of $n\in\mathbb{N}$. Recall from Remark 2.34 that the spatial derivative $v$ of a GCSF satisfies the equation
(5.1) $\partial_{t}v=\partial_{x}^{2}\!\left(\tan^{-1}(v)\!\right)=\partial_{x}\!\left(\frac{1}{1+v^{2}}\partial_{x}v\right).$
@@
(5.3) $\|u^{n}\|_{\mathrm{L}^{\infty}(J\times[\sigma,\tau])}\leq\|U^{n}\|_{\mathrm{L}^{\infty}(J\times[\sigma,\tau])}\leq C$
@@
(5.4) $u^{r,y}(x,t)=\frac{1}{r}u(y+rx,r^{2}t),$
which starts from $u^{r,y}_{0}(x)=r^{-1}u_{0}(y+rx)$, cf. (3.1). We can readily check that
(5.5) $\|u_{0}^{r,y}\|_{\mathrm{L}^{1}(J_{1})}$ $=\frac{1}{r^{2}}\|u_{0}\|_{\mathrm{L}^{1}(J_{r}(y)\!)},$
@@
(5.6) $\left|u^{r,y}\left(0,\frac{t}{r^{2}}\right)\right|$ $=\frac{1}{r}|u(y,t)|,$
@@
(5.7) $|u(y,t)|\leq Cr+\frac{1}{2r}\|u_{0}\|_{\mathrm{L}^{1}(J_{r}(y)\!)}+\frac{\pi t}{2r}$
@@
Next, suppose that $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{2r}(x)\times[0,\infty)\!)$ is a GCSF starting from $u_{0}:=u(0)\in\mathrm{L}^{1}(J_{2r}(x)\!)$ with $S\geq\|u_{0}\|_{\mathrm{L}^{1}(J_{2r}(x)\!)}$. For any $y\in J_{r}(x)$ and $T>S$, we can apply the rescaled estimate (5.7) to $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{r}(y)\times[0,T])$ to get that
(5.9) $|u(y,t)|$ $\leq Cr+\frac{1}{2r}\|u_{0}\|_{\mathrm{L}^{1}(J_{r}(y)\!)}+\frac{\pi t}{2r}$
$\leq Cr+\frac{S}{2r}+\frac{\pi T}{2r}$
@@
(5.11) $\|u\|_{\mathrm{L}^{\infty}(J_{r}(x)\times[S,T])}\leq C(r,S,T).$
We will apply our interior area-to-height estimate (Theorem 1.8) in the form (5.11).
@@
(5.13) $\int_{y\in J_{2r_{x}}(x)}U^{n}_{0}(y)\,\mathrm{d}y=\|U^{n}_{0}\|_{\mathrm{L}^{1}(J_{2r_{x}}(x)\!)}\to|\nu|(J_{2r_{x}}(x)\!)\leq\frac{\sigma}{2}$
@@
(5.15) $\|U^{n}\|_{\mathrm{L}^{\infty}(J_{r_{x}}(x)\times[\sigma,\tau])}\leq C(r_{x},\sigma,\tau)$
@@ 5.2 Convergence to the initial datum
We handle the weak convergence first. As in the analogous case of the two–dimensional RF ([TY24a]), we shall do this via an area estimate.
@@
(5.18) $\left|\frac{\mathrm{d}}{\mathrm{d}t}\int_{y\in J}\varphi(y)u(y,\cdot\,)\,\mathrm{d}y\,\right|\leq C(\varphi)=\frac{\pi}{2}\int_{y\in J}|\partial_{x}\varphi(y)|\,\mathrm{d}y.$
Proof.
@@
[…]
@@
(5.20) $\left|\,\int_{y\in\mathbb{R}}\varphi(y)u^{n}(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\varphi(y)u^{n}_{0}(y)\,\mathrm{d}y\,\right|\leq C(\varphi)t_{m}.$
@@
(5.21) $\left|\,\int_{y\in\mathbb{R}}\varphi(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\varphi(y)\,\mathrm{d}\nu(y)\,\right|\leq C(\varphi)t_{m}.$
@@
(5.23) $\|U(t_{m})\|_{\mathrm{L}^{1}(J)}\leq|\nu|(J)+\pi t_{m}.$
For a general $\varphi\in\mathrm{C}^{0}_{\mathrm{loc}}(\mathbb{R})$, we approximate by a smooth $\widetilde{\varphi}\in\mathrm{C}^{\infty}_{\mathrm{c}}(J)$ where $\operatorname{supp}(\varphi)\Subset J\Subset\mathbb{R}$. For see that
(5.24) $\left|\,\int_{y\in\mathbb{R}}\varphi(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\varphi(y)\,\mathrm{d}\nu(y)\right|$
$\leq\left|\,\int_{y\in J}(\varphi(y)-\widetilde{\varphi}(y)\!)u(y,t_{m})\,\mathrm{d}y\,\right|+\left|\,\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)\,\mathrm{d}\nu(y)\right|$
$\qquad{}+\left|\,\int_{y\in J}(\widetilde{\varphi}(y)-\varphi(y)\!)\,\mathrm{d}\nu(y)\right|$
$\leq\|\varphi-\widetilde{\varphi}\|_{\mathrm{L}^{\infty}(J)}\big{(}\|u(t_{m})\|_{\mathrm{L}^{1}(J)}+|\nu|(J)\big{)}+\left|\,\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)\,\mathrm{d}\nu(y)\right|.\!$
@@
The strong convergence (1.15) is a modification of part of [CK20, Theorem A]; we recall their argument for completeness. The key tool is the following $\mathrm{L}^{p}_{\mathrm{loc}}$–separation estimate.
@@
(5.26) $\|(u_{1}(t)-u_{2}(t)\!)\chi\|_{\mathrm{L}^{p}(J_{R})}\leq\|(u_{1}(s)-u_{2}(s)\!)\chi\|_{\mathrm{L}^{p}(J_{R})}+\frac{2\pi(1+\delta)p}{(R-r)^{1-\frac{1}{p}}}(t-s)$
for all $s\leq t\in(0,T)$.
@@
(5.29) $\|u_{1}(t)-u_{2}(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\|u_{1}(s)-u_{2}(s)\|_{\mathrm{L}^{p}(J_{R})}+C_{p,R-r}(t-s)$
for all $s<t\in(0,T)$.
Let $\mathcal{K}\subset\Omega$ be compact; then as $\Omega\subset\mathbb{R}$ is open, and therefore a countable union of disjoint open intervals, it follows that $\mathcal{K}$ is contained within a finite union of open intervals. Thus, it suffices to prove that $u(t)\to u_{0}\in\mathrm{L}^{p}(J)$ for a single open interval $J\Subset\Omega$. By translation, we may suppose that $J=J_{r}\Subset J_{R}\subseteq\Omega$.
Let $\varepsilon>0$. We first choose $m\in\mathbb{N}$ so large that
(5.30) $\|u^{m}_{0}-u_{0}\|_{\mathrm{L}^{p}(J_{R})}\leq\frac{\varepsilon}{4}.$
With this $m$ fixed, we then choose $t_{\varepsilon}>0$ so small that
(5.31) $\|u^{m}(t)-u(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\frac{\varepsilon}{2}\qquad\text{and }\qquad\|u^{m}(t)-u^{m}_{0}\|_{\mathrm{L}^{p}(J_{r})}\leq\frac{\varepsilon}{4}$
for all $0<t<t_{\varepsilon}$, using (5.29) and the locally smooth convergence $u^{m}(t)\to u^{m}_{0}$ as $t\searrow 0$ respectively. Taken altogether, we get that
(5.32) $\|u(t)-u_{0}\|_{\mathrm{L}^{p}(J_{r})}$ $\leq\|u(t)-u^{m}(t)\|_{\mathrm{L}^{p}(J_{r})}+\|u^{m}(t)-u^{m}_{0}\|_{\mathrm{L}^{p}(J_{r})}+\|u^{m}_{0}-u_{0}\|_{\mathrm{L}^{p}(J_{r})}$
@@ 5.3 Some comments on the work of Chou & Kwong
First, we clarify that if the initial datum $\nu=u_{0}\mathscr{L}^{1}\in\mathcal{M}_{*}(\mathbb{R})$ has density $u_{0}\in\mathrm{L}^{p}(\mathbb{R})$ for some $1\leq p<\infty$, then the solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ constructed in Theorem 1.13 converges in $\mathrm{L}^{p}(\mathbb{R})$.
Sketch proof of (1.16).
@@
(5.33) $\|u(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\|u_{0}\|_{\mathrm{L}^{p}(J_{R})}+C_{p,R-r}t.$
Letting $r=R-1\nearrow\infty$ gives us that
(5.34) $\|u(t)\|_{\mathrm{L}^{p}(\mathbb{R})}\leq\|u_{0}\|_{\mathrm{L}^{p}(\mathbb{R})}+C_{p,1}t,$
@@
(5.35) $\limsup_{t\searrow 0}\|u(t)\|_{\mathrm{L}^{p}(\mathbb{R})}\leq\|u_{0}\|_{\mathrm{L}^{p}(\mathbb{R})}.$
The following measure-theoretic lemma is straightforward to prove:
@@
Let $(f_{n})_{n\in\mathbb{N}}\subset\mathrm{L}^{p}(\mathbb{R})$ be a sequence of functions for some fixed $1\leq p<\infty$, and let $f\in\mathrm{L}^{p}(\mathbb{R})$. Suppose that
(5.37) $f_{n}\to f\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\qquad\text{ and }\qquad\limsup_{n\nearrow\infty}\|f_{n}\|_{\mathrm{L}^{p}(\mathbb{R})}\leq\|f\|_{\mathrm{L}^{p}(\mathbb{R})}.$
@@
(5.38) $f_{n}\to f\in\mathrm{L}^{p}(\mathbb{R}).$
Applying Lemma 5.36 to $f_{n}=u(t_{n})$ and $f=u_{0}$ for an arbitrary sequence $(t_{n})_{n\in\mathbb{N}}\subset(0,\infty)$ converging to zero, keeping in mind (1.15) and (5.35), gives the desired global convergence (1.16). ∎
Second, we highlight that Proposition 5.25 implies the following $\mathrm{L}^{p}_{\mathrm{loc}}$–continuity property for the GCSF when $1<p<\infty$.
Theorem 5.39 ($\mathrm{L}^{p>1}_{\mathrm{loc}}(\mathbb{R})$-to-$\mathrm{C}^{0}_{\mathrm{loc}}([0,\infty);\mathrm{L}^{p>1}_{\mathrm{loc}}(\mathbb{R})\!)$–continuity).
Let $1<p<\infty$, and let $u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$. Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ be any GCSF for which $u(t)\to u(0):=u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$ as $t\searrow 0$. Then $u\in\mathrm{C}^{0}_{\mathrm{loc}}([0,\infty);\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\!)$, and the mapping $\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\ni u_{0}\to u\in\mathrm{C}^{0}_{\mathrm{loc}}([0,\infty);\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\!)$ is well-defined and continuous; in particular, for each $u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$, the associated solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ is unique.
Proof.
@@
The rest of the theorem follows from Proposition 5.25. Fix $J_{r}\Subset\mathbb{R}$ and $[a,b]\Subset[0,\infty)$, and let $\varepsilon>0$ be arbitrary. Choose $R>0$ so large that
(5.40) $r<\frac{R}{2}\qquad\text{ and }\qquad\frac{2^{3-\frac{1}{p}}\pi pt}{R^{1-\frac{1}{p}}}\leq\frac{8\pi pb}{R^{1-\frac{1}{p}}}\leq\frac{\varepsilon}{2},$
@@
(5.41) $\|u_{0}-\widetilde{u}_{0}\|_{\mathrm{L}^{p}(J_{R})}\leq\frac{\varepsilon}{2},$
@@
(5.42) $\|u(t)-\widetilde{u}(t)\|_{\mathrm{L}^{p}(J_{r})}$ $\leq\|u(t)-\widetilde{u}(t)\|_{\mathrm{L}^{p}(J_{R/2})}$

### v2

@@ 5.1 Constructing the flow
We start with a sequence $(u^{n}_{0})_{n\in\mathbb{N}}\subset\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R})$ of smooth and compactly supported functions for which $u^{n}_{0}\mathscr{L}^{1}\rightharpoonup\nu\in\mathcal{M}(\mathbb{R})$ and $u^{n}_{0}\to u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\Omega)$, and with another sequence $(U^{n}_{0})_{n\in\mathbb{N}}\subset\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R})$ for which $U_{0}^{n}\mathscr{L}^{1}\rightharpoonup|\nu|\in\mathcal{M}_{*}^{+}(\mathbb{R})$ and $U_{0}^{n}\to|u_{0}|\in\mathrm{L}^{p}_{\mathrm{loc}}(\Omega)$, and which is chosen to satisfy $|u^{n}_{0}|\leq U^{n}_{0}$ on $\mathbb{R}$ for each $n\in\mathbb{N}$. We can construct such sequences by cutting-off and mollifying the positive and negative parts of $\nu\in\mathcal{M}_{*}(\mathbb{R})$ simultaneously, and then recombining the mollifications (see e.g. [Rud91, Ch. 6]).
Let $u^{n}\in\mathrm{C}^{\infty}(\mathbb{R}\times[0,\infty)\!)$ be the Ecker–Huisken flow of $u^{n}_{0}$, and let $U^{n}\in\mathrm{C}^{\infty}(\mathbb{R}\times[0,\infty);\mathbb{R}_{\geq 0})$ be the Ecker–Huisken flow of $U^{n}_{0}$. The Avoidance Principle (Theorem A.7) applied to $\pm u^{n}$ and $U^{n}+\eta$ for $\eta\searrow 0$ tells us that $|u^{n}(t)|\leq U^{n}(t)$ on $\mathbb{R}$ for all $t\in[0,\infty)$.
We wish to take a locally smooth limit of the flows $(u^{n})$ on $\mathbb{R}\times(0,\infty)$; for technical reasons in §5.2, we will also need to take a limit of the flows $(U^{n})$. By the Arzelà–Ascoli Theorem, it suffices to prove $\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$–bounds on the sequences of GCSFs which are independent of $n\in\mathbb{N}$. Recall from (2.35) that the spatial derivative $v=\partial_{x}u$ of a GCSF $u$ satisfies the equation
$\partial_{t}v=\partial_{x}^{2}[\tan^{-1}(v)]=\partial_{x}\left[\frac{1}{1+v^{2}}\partial_{x}v\right].$ (5.1)
@@
$\|u^{n}\|_{\mathrm{L}^{\infty}(J\times[\sigma,\tau])}\leq\|U^{n}\|_{\mathrm{L}^{\infty}(J\times[\sigma,\tau])}\leq C$ (5.3)
@@
$u^{r,y}(x,t):=\frac{1}{r}u(y+rx,r^{2}t),$ (5.4)
which starts from $u^{r,y}_{0}(x):=r^{-1}u_{0}(y+rx)$, cf. (3.1). We can readily check that
$\|u_{0}^{r,y}\|_{\mathrm{L}^{1}(J_{1})}=\frac{1}{r^{2}}\|u_{0}\|_{\mathrm{L}^{1}(J_{r}(y)\!)},$ (5.5)
@@
$\left|u^{r,y}\left(0,\frac{t}{r^{2}}\right)\right|=\frac{1}{r}|u(y,t)|,$ (5.6)
@@
$|u(y,t)|\leq Cr+\frac{1}{2r}\|u_{0}\|_{\mathrm{L}^{1}(J_{r}(y)\!)}+\frac{\pi t}{2r}$ (5.7)
@@
Next, suppose that $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{2r}(x)\times[0,\infty)\!)$ is a GCSF starting from $u_{0}:=u(0)\in\mathrm{L}^{1}(J_{2r}(x)\!)$. For any $y\in J_{r}(x)$ and $\|u_{0}\|_{\mathrm{L}^{1}(J_{2r}(x)\!)}\leq S\leq T$, we can apply the rescaled estimate (5.7) to $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{r}(y)\times[0,T])$ to get that
$|u(y,t)|\leq Cr+\frac{1}{2r}\|u_{0}\|_{\mathrm{L}^{1}(J_{r}(y)\!)}+\frac{\pi t}{2r}\leq Cr+\frac{S}{2r}+\frac{\pi T}{2r}$ (5.9)
@@
$\|u\|_{\mathrm{L}^{\infty}(J_{r}(x)\times[S,T])}\leq C(r,S,T)$ (5.11)
for all $\|u_{0}\|_{\mathrm{L}^{1}(J_{2r}(x)\!)}\leq S\leq T$. We will apply our interior area-to-height estimate (Theorem 1.8) in the form (5.11).
@@
$\|U^{n}_{0}\|_{\mathrm{L}^{1}(J_{2r_{x}}(x)\!)}=\int_{y\in J_{2r_{x}}(x)}U^{n}_{0}(y)\,\mathrm{d}y\to\int_{y\in J_{2r_{x}}(x)}\,\mathrm{d}|\nu|(y)=|\nu|(J_{2r_{x}}(x)\!)\leq\frac{\sigma}{2}$ (5.13)
@@
$\|U^{n}\|_{\mathrm{L}^{\infty}(J_{r_{x}}(x)\times[\sigma,\tau])}\leq C(r_{x},\sigma,\tau)$ (5.15)
@@ 5.2 Convergence to the initial datum
We handle the weak convergence first. As in the analogous case of the two–dimensional RF in [TY24a], we shall do this via an area estimate.
@@
$\left|\frac{\mathrm{d}}{\mathrm{d}t}\int_{y\in J}\varphi(y)u(y,\cdot\,)\,\mathrm{d}y\,\right|\leq C(\varphi):=\frac{\pi}{2}\int_{y\in J}|\partial_{x}\varphi(y)|\,\mathrm{d}y.$ (5.18)
?proofname?.
@@
$\begin{aligned} \left|\,\left.\frac{\mathrm{d}}{\mathrm{d}t}\right.\int_{y\in J}\varphi(y)u(y,\cdot)\,\mathrm{d}y\,\right|&=\left|\,\int_{y\in J}\varphi(y)\cdot\partial_{x}[\tan^{-1}(v)](y,s)\,\mathrm{d}y\,\right|\\ &=\left|\,\int_{y\in J}\partial_{x}\varphi(y)\cdot[\tan^{-1}(v)](y,s)\,\mathrm{d}y\,\right|\\ &\leq\frac{\pi}{2}\int_{y\in J}|\partial_{x}\varphi(y)|\,\mathrm{d}y.\end{aligned}\qed$ (5.19)
@@
$\left|\,\int_{y\in\mathbb{R}}\varphi(y)u^{n}(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\varphi(y)u^{n}_{0}(y)\,\mathrm{d}y\,\right|\leq C(\varphi)t_{m}.$ (5.20)
@@
$\left|\,\int_{y\in\mathbb{R}}\varphi(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\varphi(y)\,\mathrm{d}\nu(y)\,\right|\leq C(\varphi)t_{m}.$ (5.21)
@@
$\|U(t_{m})\|_{\mathrm{L}^{1}(J)}\leq|\nu|(J)+\pi t_{m}.$ (5.23)
For a general $\varphi\in\mathrm{C}^{0}_{\mathrm{c}}(\mathbb{R})$, we approximate by a smooth $\widetilde{\varphi}\in\mathrm{C}^{\infty}_{\mathrm{c}}(J)$ where $\operatorname{supp}(\varphi)\Subset J\Subset\mathbb{R}$. For see that
$\left|\,\int_{y\in\mathbb{R}}\varphi(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\varphi(y)\,\mathrm{d}\nu(y)\right|$ (5.24)
$\qquad{}\leq\left|\,\int_{y\in J}[\varphi(y)-\widetilde{\varphi}(y)]u(y,t_{m})\,\mathrm{d}y\,\right|$
$\qquad\qquad{}+\left|\,\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)\,\mathrm{d}\nu(y)\right|$
$\qquad\qquad{}+\left|\,\int_{y\in J}[\widetilde{\varphi}(y)-\varphi(y)]\,\mathrm{d}\nu(y)\right|$
$\qquad\leq\|\varphi-\widetilde{\varphi}\|_{\mathrm{L}^{\infty}(J)}\big[\|u(t_{m})\|_{\mathrm{L}^{1}(J)}+|\nu|(J)\big]$
$\qquad\qquad{}+\left|\,\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)u(y,t_{m})\,\mathrm{d}y-\int_{y\in\mathbb{R}}\widetilde{\varphi}(y)\,\mathrm{d}\nu(y)\right|.\!$
@@
The strong convergence (1.16) is a modification of part of [CK20, Theorem A]; we recall their argument for completeness. The key tool is the following $\mathrm{L}^{p}_{\mathrm{loc}}$–separation estimate:
@@
$\|(u_{1}(t)-u_{2}(t)\!)\chi\|_{\mathrm{L}^{p}(J_{R})}\leq\|(u_{1}(s)-u_{2}(s)\!)\chi\|_{\mathrm{L}^{p}(J_{R})}+\frac{2\pi(1+\delta)p}{(R-r)^{1-\frac{1}{p}}}(t-s)$ (5.26)
for all $0<s\leq t<T$.
@@
$\|u_{1}(t)-u_{2}(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\|u_{1}(s)-u_{2}(s)\|_{\mathrm{L}^{p}(J_{R})}+C_{p,R-r}(t-s)$ (5.29)
for all $0<s\leq t<T$.
Let $\mathcal{K}\Subset\Omega$ be compact; then as $\Omega\subseteq\mathbb{R}$ is open, and therefore a countable union of disjoint open intervals, it follows that $\mathcal{K}$ is contained within a finite union of open intervals. Thus, it suffices to prove that $u(t)\to u_{0}\in\mathrm{L}^{p}(J)$ for a single open interval $J\Subset\Omega$. By translation, we may suppose that $J=J_{r}\Subset J_{R}\subseteq\Omega$.
Let $\varepsilon>0$. We first choose $n\in\mathbb{N}$ so large that
$\|u^{n}_{0}-u_{0}\|_{\mathrm{L}^{p}(J_{R})}\leq\frac{\varepsilon}{4}.$ (5.30)
With this $n$ fixed, we then choose $t_{\varepsilon}>0$ so small that
$\|u^{n}(t)-u(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\frac{\varepsilon}{2}\qquad\text{ and }\qquad\|u^{n}(t)-u^{n}_{0}\|_{\mathrm{L}^{p}(J_{r})}\leq\frac{\varepsilon}{4}$ (5.31)
for all $0<t<t_{\varepsilon}$, using (5.29) and the locally smooth convergence $u^{n}(t)\to u^{n}_{0}$ as $t\searrow 0$ respectively. Taken altogether, we get that
$\|u(t)-u_{0}\|_{\mathrm{L}^{p}(J_{r})}$ $\leq\|u(t)-u^{m}(t)\|_{\mathrm{L}^{p}(J_{r})}+\|u^{m}(t)-u^{m}_{0}\|_{\mathrm{L}^{p}(J_{r})}+\|u^{m}_{0}-u_{0}\|_{\mathrm{L}^{p}(J_{r})}$ (5.32)
@@ 5.3 Some comments on the work of Chou & Kwong
First, we clarify that if the initial datum $\nu=u_{0}\mathscr{L}^{1}\in\mathcal{M}_{*}(\mathbb{R})$ has density $u_{0}\in\mathrm{L}^{p}(\mathbb{R})$ for some $1\leq p<\infty$, then the solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ constructed in Theorem 1.14 attains its initial datum in $\mathrm{L}^{p}(\mathbb{R})$.
Proof of (1.17).
@@
$\|u(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\|u_{0}\|_{\mathrm{L}^{p}(J_{R})}+C_{p,R-r}t.$ (5.33)
Letting $r:=R-1\nearrow\infty$ gives us that
$\|u(t)\|_{\mathrm{L}^{p}(\mathbb{R})}\leq\|u_{0}\|_{\mathrm{L}^{p}(\mathbb{R})}+C_{p,1}t,$ (5.34)
@@
$\limsup_{t\searrow 0}\|u(t)\|_{\mathrm{L}^{p}(\mathbb{R})}\leq\|u_{0}\|_{\mathrm{L}^{p}(\mathbb{R})}.$ (5.35)
The following measure-theoretic lemma is straightforward to prove (see e.g. [Sob24, §4.3]):
@@
Let $(f_{m})_{m\in\mathbb{N}}\subset\mathrm{L}^{p}(\mathbb{R})$ be a sequence of functions for some fixed $1\leq p<\infty$, and let $f\in\mathrm{L}^{p}(\mathbb{R})$. Suppose that
$f_{m}\to f\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\qquad\text{ and }\qquad\limsup_{m\nearrow\infty}\|f_{m}\|_{\mathrm{L}^{p}(\mathbb{R})}\leq\|f\|_{\mathrm{L}^{p}(\mathbb{R})}.$
@@
$f_{m}\to f\in\mathrm{L}^{p}(\mathbb{R}).$ (5.37)
Applying Lemma 5.36 to $f_{m}:=u(t_{m})$ and $f:=u_{0}$ for an arbitrary sequence $(t_{m})_{m\in\mathbb{N}}\subset(0,\infty)$ converging to zero, keeping in mind (1.16) and (5.35), gives the desired global convergence (1.17). ∎
Second, we highlight that Proposition 5.25 implies the following continuity property for the GCSF in $\mathrm{L}^{p}_{\mathrm{loc}}$ for $1<p<\infty$:
Theorem 5.38 (Continuous dependence on $\mathrm{L}^{p}_{\mathrm{loc}}$–initial data when $p>1$).
Let $1<p<\infty$, and let $u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$. Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ be any GCSF for which $u(t)\to u(0):=u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$ as $t\searrow 0$. Then $u\in\mathrm{C}^{0}_{\mathrm{loc}}([0,\infty);\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\!)$, and the mapping $\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\ni u_{0}\mapsto u\in\mathrm{C}^{0}_{\mathrm{loc}}([0,\infty);\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})\!)$ is well-defined and continuous; in particular, for each $u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$, the associated solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ is unique.
?proofname?.
@@
The rest of the theorem follows from Proposition 5.25. Fix $J_{r}\Subset\mathbb{R}$ and $a,b]\Subset[0,\infty)$, and let $\varepsilon>0$ be arbitrary. Choose $R>0$ so large that
$r<\frac{R}{2}\qquad\text{ and }\qquad\frac{2^{3-\frac{1}{p}}\pi pt}{R^{1-\frac{1}{p}}}\leq\frac{8\pi pb}{R^{1-\frac{1}{p}}}\leq\frac{\varepsilon}{2},$ (5.39)
@@
$\|u_{0}-\widetilde{u}_{0}\|_{\mathrm{L}^{p}(J_{R})}\leq\frac{\varepsilon}{2},$ (5.40)
@@
$\|u(t)-\widetilde{u}(t)\|_{\mathrm{L}^{p}(J_{r})}$ $\leq\|u(t)-\widetilde{u}(t)\|_{\mathrm{L}^{p}(J_{R/2})}$ (5.41)